OpenStudy (anonymous):

How to solve this 1/a + 1/b = b/a you need to find a?

OpenStudy (anonymous):

Somebody help me

OpenStudy (whpalmer4):

\[\frac{1}{a}+\frac{1}{b} = \frac{b}{a}\]To solve for \(a\), you need to either get a common denominator or clear out the fractions. Do you know how to do either of those things?